# Sharing Matters: Analysing Neurons Across Languages and Tasks in LLMs

### v2

@@ Abstract
Large language models (LLMs) have revolutionized the field of natural language processing (NLP), and recent studies have aimed to understand their underlying mechanisms. However, most of this research is conducted within a monolingual setting, primarily focusing on English. Few studies attempt to explore the internal workings of LLMs in multilingual settings. In this study, we aim to fill the research gap by examining how neuron activation is shared across tasks and languages. We classify neurons into four distinct categories based on their responses to a specific input across different languages: all-shared, partial-shared, specific, and non-activated. This categorization is combined with a study of neuron attribution, i.e. the importance of a neuron w.r.t an output. Our analysis reveals the following insights: (i) the patterns of neuron sharing are significantly affected by the characteristics of tasks and examples; (ii) neuron sharing does not fully correspond with language similarity; (iii) shared neurons play a vital role in generating responses, especially those shared across all languages. These findings shed light on the internal workings of multilingual LLMs and pave the way to the future research. We will release the code to foster research in this area.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in recent studies, excelling in both understanding and generating text across various languages (OpenAI, 2023; Zhang et al., 2023; Zhao et al., 2024). Despite their proven effectiveness, the intricate mechanisms underlying their processing remain largely opaque. This opacity has given rise to a growing field of research aimed at interpreting the internal workings of the Transformer architecture (Elhage et al., 2021; Yu et al., 2023). To enhance interpretability and investigate specific aspects of model behavior, researchers have increasingly focused on the components of these models. Recent studies have explored the role of Feed-Forward Networks (FFNs) within LLMs, proposing that these components function as key-value memories for storing factual and linguistic knowledge (Geva et al., 2020, 2022; Ferrando et al., 2023). While these studies have analyzed neuron behaviors based on activation states in monolingual settings, there remains a significant gap in our understanding of how neurons behave in multilingual contexts.
@@
Significant discrepancies have been observed in neuron activation patterns across languages, despite identical semantic outputs, as highlighted in Figure 1. This phenomenon raises critical questions about how LLMs manage multilingual representations at the neuronal level. In a similar line of reasoning, Bhattacharya and Bojar (2023) categorized neurons into two coarse-grained groups: language-agnostic (shared across languages) and language-specific (unique to a language). However this categorization oversimplifies the complexity observed in cross-lingual studies, where neuron overlap varies significantly between languages (Stanczak et al., 2022; Zhao et al., 2023; Liu et al., 2024). Additionally, most research has been confined to single-task analyses, overlooking how neuron types might shift across diverse tasks (Bhattacharya and Bojar, 2023; Tang et al., 2024; Tan et al., 2024). This underscores the need for a more nuanced, fine-grained classification method to enhance our understanding of the multifaceted roles of neurons in multilingual LLMs.
@@
To bridge this gap, our research introduces a fine-grained classification of neurons, enabling a detailed exploration of their roles across languages and tasks. We systematically analyse neuron behaviors across three distinct tasks in ten languages, including reasoning, fact probing, and question answering tasks, using various diverse model backbones: BLOOMZ-7b, LLaMA2-7b-chat, BLOOM-7b, XGLM, and mT0. For a specific input and its corresponding translations in 10 languages, we reformulate the activation state of neurons to four distinctive types to represent multilingual neuron roles (see Section 3). All-shared neurons are neurons that remain active for all the translations regardless of language. Partial-shared neurons are activated only for translations in certain languages. Specific neurons are activated exclusively for translations in one language. Non-activated neurons are not activated at all for any the input’s translations. We then use the ratio of neurons in each type to analyze the activation pattern across various tasks. Furthermore, we evaluate the importance of each type of neurons in predicting answers using contribution scores and effective scores (Geva et al., 2022; Voita et al., 2023). Finally, we confirm our analysis through experiments on downstream tasks.
@@
We provide substantial empirical evidence detailing neuron activation patterns and attributions in this study, leading to several significant findings. Here are the main takeaways:
Neuron activation patterns are considerably task-specific and example-specific. We observe the ratio of four types of neurons is related to the task (see Section 5.1), where the neuron activation pattern of XNLI is significantly different from that of Fact Probing. We also demonstrate that the neuron types commonly vary with different test examples (see Section 5.2).
@@
Similar languages do not always lead to high degree of neuron sharing. Our empirical results show that the languages from the same language family, such as German and French, do not exhibit higher degree of neuron sharing, compared to the languages from distinct language families, such as German and Chinese (see Section 5.3).
Shared neurons are crucial in generating responses. Both the contribution scores and the effective scores highlight the significance of all-shared neurons (see Section 6). Our experiments further reveal that disabling the all-shared neurons can lead to a substantial decrease (up to 91.6%) in model performance on downstream tasks (see Section 7).
@@ 2 Related Work
For multilingual neuron analysis, Bhattacharya and Bojar (2023) explored the neuron sharing between two languages. Tang et al. (2024); Tan et al. (2024); Liu et al. (2024); Kojima et al. (2024) classified neurons in an FFN block to language-specific and language-agnostic based on predefined threshold, assuming that a neuron’s classified type remained consistent across different examples. However, the broad classification into two groups is inadequate for detailed multilingual analysis. Additionally, these studies classified neurons based on the single task (Tan et al., 2024; Liu et al., 2024), without considering the potential adaptation of neurons under various languages and semantics brought forth by inputs from various multilingual tasks. We investigate neurons’ behaviors across multiple languages and tasks to this end.
3 Neuron Classification
In this section, we provide a detailed description of the 4-way neuron classification that we propose. We begin with some background concerning neurons in the FFN block (Section 3.1). Following this, we define the four types of neurons (Section 3.2), and explain how we collect activation patterns by averaging the ratio of each neuron type across the entire test set (Section 3.3).
@@ 3.1 Neurons in FFN Blocks
A neuron inside the FFNs is defined as a linear transformation of an input representation followed by a non-linear activation (Tang et al., 2024). Every feed-forward block at layer involves two linear transformations separated by a point-wise activation function. Biases are omitted for the sake of clarity:
where , are linear parameter matrices, and is a non-linear activation function, where rows in and columns in are viewed as -dimensional keys and values , respectively. We define the behavior of a neuron to be the output of the function immediately after the elementwise nonlinearity. is the count of neurons. And the output of neurons determines the weighting of the corresponding values in .
@@
When such a neuron is activated , so it updates the residual stream by pulling out the corresponding value . If a neuron is inactive, the activation value is .
3.2 Definition of Four Types of Neurons
In this work, we categorize the neurons into four types based on their activation values and detail the neuron classification in this section. To ablate the impact of semantic discrepancies across languages, the datasets used in this work are initially in English and then translated into foreign languages (see Section 4.1), so we can formulate the -th example as , where indicates the -th language and is the total number of languages.
Given the -th example , the set of all-shared neurons at the -th layer can be defined as:
where is the set of all the neurons at the -th layer and is the -th neuron in . Furthermore, the non-activated neurons is the set of neurons whose activation value less than or equal to zero in all languages, as follows:
@@ 3.2 Definition of Four Types of Neurons
Note that, unlike Bhattacharya and Bojar (2023); Tang et al. (2024) who consider the sub-word activation states and thus capture incomplete semantics, we only examine the activation state of the last token, as that is when the LLM performs the prediction task.
3.3 Collecting Neuron Activation Patterns
In this section, we explain how to quantify neuron activation patterns. Specifically, we measure the percentage of each type of neuron relative to the total number of neurons. Given the -th test instance, the percentage of each neuron type at the -th layer can be defined as follows:
where denotes the number of elements in the set. Consequently, the aggregated neuron activation pattern for one dataset contain instances at the -th layer can be defined as:
@@ 4.1 Multilingual Tasks
We perform analysis on neurons in FFN blocks of various LLMs, harnessing their multilingual capabilities in three diverse tasks which consist of multilingual parallel sentences, including XNLI (Conneau et al., 2018), Fact Probing (Fierro and Søgaard, 2022), and Cross-lingual Knowledge Editing (KE) (Wang et al., 2023). For the Cross-lingual KE, we analyze the LLMs in two setups, including EN (Edit) ALL (Test) and ALL (Edit) EN (Test). More details are described in Appendix A.2.
@@
These tasks cover 10 diverse languages, including English (en), German (de), Spanish (es), French (fr), Portuguese (pt), Russian (ru), Thai (th), Turkish (tr), Vietnamese (vi), and Chinese (zh). Prompts are detailed in Appendix A.3.
4.2 Base LLMs
@@
We mainly analyze the behaviors and attributions of neurons in an instruction-finetuned model BLOOMZ-7b (Muennighoff et al., 2023). We also include the analysis of other decoder-only models: BLOOM-7b (Scao et al., 2022), LLAMA2-7b-chat (Touvron et al., 2023), XGLM (Lin et al., 2022), and an encoder-decoder model mT0 (Muennighoff et al., 2023) in the Appendix A.4.1.
5 Behaviors of Four Types of Neurons
In this section, we present our in-depth analysis with regard to the neuron activation across task (Section 5.1), examples within the same task (Section 5.2), and language (Section 5.3).
For each task we use parallel test texts in ten languages as input, and record the type of each neuron. Subsequently, we calculate and compare the percentage of the four types of neurons compared to the total neurons. We use BLOOMZ-7b as the backbone LLM to investigate the behaviors of four types of neurons, and the results of three tasks are shown in Figure 3 respectively.
5.1 Neuron Activation Across Tasks
The neuron activation pattern is strongly related to the task performed.
As shown in Figure 3, non-activated neurons are more prevalent overall than other types. However, the pattern observed in the Fact Probing task differs significantly from the other two tasks. In the Fact Probing task, there are more partial-shared neurons and specific neurons, whereas the other tasks involve far more all-shared neurons. In LABEL:bloom-llama-3tasks, the activation patterns of the BLOOM-7b and LLaMA2-7b-chat LLMs are presented, highlighting variability across models while affirming the consistent influence of the task on neuron activation patterns. Additionally, a comparison between BLOOM and BLOOMZ LLMs indicates that instruction fine-tuning in BLOOMZ appears to foster a higher proportion of all-shared neurons, suggesting better alignment across multilingual representations.
5.2 Neuron Activation Across Examples
The type of a neuron varies across examples within the same task.
As shown in Figure 3, we have analysed the percentage of each type of neurons across tasks. In this experiment, we now investigate how neurons behave across all examples within one task, aiming to determine whether neurons categorized as a specific type retain their classification across different semantics. Take the Fact Probing task as an example, the results are shown in Figure 5 and in Appendix A.4.2 (for other tasks). To our surprise, almost no neuron (identified by its index) retains its type (in our four-ways classification) across all examples. This finding contradicts prior research that suggested neuronal classification remains unchanged amidst variation in input semantics. Our study thus underscores that the classified type of a neuron is sensitive to the semantics of the encountered inputs, even within the same task.
Non-activated neurons do not stay inactive across examples.
In Figure 5 we observe that less than 10% of non-activated neurons remain inactive across all examples. This small proportion of persistently inactive neurons reflects the distributed nature of knowledge representation in LLMs. That is why we should execute caution when pruning neurons, as we may damage the overall performance of an LLM.
Non-activated neurons do not stay inactive across all tokens in an input.
Expanding our focus, we examine whether non-activated neurons remain inactive across all tokens of an input, differing from our prior analysis which concentrated only on the last token. Specifically, for each input sequence with tokens , we record non-activated neurons of each token, w.r.t the intersection of index. A neuron is counted when it stays inactive for every token of one test input sequence. As shown in Figure 6, fewer than 0.8% non-activated neurons consistently stay inactive for each token of the input, which further highlights the variability in neuron activity.
5.3 Neuron Activation Across Languages
Neuron sharing does not completely align with language similarity.
In exploring the relationship between language similarity and neuron sharing, we specifically analyse the proportion of shared neurons between German and other languages within the category of partial-shared neurons. Additionally, we examine the ratio of specific neurons for each language. These analyses are conducted within the Fact Probing task. As shown in Figure 7, similar languages do not share more neurons; for instance, German and Chinese (in different language families) have a similar ratio of shared neurons to that of German and French (in the same language family) in the upper layers. Furthermore, we observe that there is no consistent pattern with regard to the ratio of specific neurons across the languages studied. We leave the further exploration on this matter to the future work. More results on XNLI task can be found in Appendix A.4.3.
More languages encourage greater neuron sharing.
We also investigate the impact of the number of languages on the percentage of four types of neurons. As illustrated in Figure 8, our study include configurations of 3, 5, 7, and 9 languages on the XNLI task. Results show a noticeable escalation in the percentage of partial-shared neurons as more languages are introduced, indicating that language configuration affects the results of model’s neuron sharing. This analysis emphasizes the potential influence of language selection on multilingual interpretability analysis.
6 Neuron Attributions
In the preceding experiments, we examined the proportions of the four types of neurons across tasks and languages. In this section, we evaluate importance of neurons in answer generation using contribution scores (Section 6.2), and assess the influence of neurons in predicting correct answers through effective scores (Section 6.3) with BLOOMZ-7b backbone LLM. We start with the fundamentals of the contribution analysis method (Section 6.1).
6.1 Contribution Score and Effective Score
In this section, we introduce two measures of quantifying the importance of neurons: Contribution Score and Effective Score.
Contribution Score
Inspired by Geva et al. (2022), in order to judge the importance of neurons in generating answers, we analyze their contributions to the output. The contribution score of the -th neuron at -th layer to an FFN output is:
which is the proportion of its weight to the sum of weights of all neurons in the FFN block. is the absolute value of activation value and is the L2-norm of value .
Effective Score
Following Geva et al. (2022) and Voita et al. (2023), the effective score of the -th neuron at -th layer can be defined as:
where is the embedding of the correct answer . A larger has a higher probability to produce the correct answer , while a negative reduces the probability in generating . In this way, we can quantify the effect of a neuron on the reference answer. We give detailed descriptions about the neuron projection in Appendix A.1.
Comparison
The contribution score quantifies how much neurons contribute to the overall generation process, regardless of the correctness of the output, while the effective score incorporates the correct answer’s embedding to measure the neurons’ impact specifically in generating accurate responses. Although both metrics assess neuronal influence, the key difference lies in their consideration of answer correctness: the contribution score focuses solely on the generation process, whereas the effective score takes into account the neurons’ role in the accuracy of the resulting answer.
6.2 The Contribution of Neuron Types
In this section, we explore the influence of each neuron type using the contribution score described in Section 6.1. We examine the proportion of the four types of neurons among the top 5% contribution score under inputs in each language. The proportions of neurons in the Cross-lingual KE task are depicted in Figure 9 (and the overall results are shown in Figure 21 in Appendix A.5.1).
All-shared neurons are the top contributing neurons to outputs.
Figure 9 depicts a stacked column chart where each segment color represents a specific neuron type’s contribution, collectively totaling 100%. It can be observed that all-shared neurons constitute the largest segment, highlighting their significant role irrespective of language input. Partial-shared neurons are the second most influential, affecting mainly the latter half of the model. It is not surprising that the specific neurons group has limited influence on Cross-lingual KE (EN (Edit) ALL (Test)) outputs as they feature in a particular language type of inputs.
We further analyze the average and total contribution scores of all neurons across four categories. As shown in Figure LABEL:bloomz-contribution_absolute-allspecific_twolangs, all-shared neurons score the highest averages, especially in upper layers, but they constitute only a minor portion of the neuron population (in Figure 3). Meanwhile, non-activated and partial-shared neurons contribute significantly to the total scores in the XNLI and fact-probing tasks respectively, predominantly due to their larger numbers. The disparity in total contributions across tasks confirms our earlier findings in Section 5.1. We leave further exploration regarding to task-specific neuron activation to the future work.
| language | neuron type | max | min | avg | var |
| en | all-shared | 1.85 | -0.94 | 0.07 | 0.36 |
| partial-shared | 0.22 | -0.16 | 1.2e-4 | 1.9e-4 | |
| specific | 0.02 | -0.02 | 2.5e-4 | 3.5e-5 | |
| non-activated | 0.04 | -0.03 | 2.1e-4 | 5.8e-6 | |
| de | all-shared | 1.03 | -0.60 | 0.02 | 0.07 |
| partial-shared | 0.13 | -0.13 | 6.7e-5 | 7.1e-5 | |
| specific | 0.07 | -0.03 | 7.5e-5 | 2.3e-5 | |
| non-activated | 0.02 | -0.01 | 7.0e-5 | 2.9e-6 |
6.3 The Effectiveness of Neuron Types
In this section, we explore the effectiveness of each neuron type using the effective score described in Section 6.1.
All-shared neurons have a greater influence on generating correct answers.
In the Cross-lingual KE (EN (Edit) ALL (Test)) task, we analyze the effectiveness scores (maximum, minimum, average, and variance) for each neuron type throughout all layers of the BLOOMZ model, as shown in Table 1. All-shared neurons display both the highest maximum and the lowest minimum scores, suggesting they are pushed (or eliminated) strongly by the activation functions. By comparison, while all-shared neurons and partial-shared neurons show a wide variance in scores (1.85 vs. -0.94 and 0.22 vs. -0.16, respectively, in English), specific neurons and non-activated neurons have much tighter score ranges ( 0.07). Additionally, we demonstrate the maximum effective scores across layers in Figure 11. It could be observed the neurons in the upper layers play a more crucial role in determining model performance.

### v3

@@ Abstract
Large language models (LLMs) have revolutionized the field of natural language processing (NLP), and recent studies have aimed to understand their underlying mechanisms. However, most of this research is conducted within a monolingual setting, primarily focusing on English. Few studies have attempted to explore the internal workings of LLMs in multilingual settings. In this study, we aim to fill this research gap by examining how neuron activation is shared across tasks and languages. We classify neurons into four distinct categories based on their responses to a specific input across different languages: all-shared, partial-shared, specific, and non-activated. Building upon this categorisation, we conduct extensive experiments on three tasks across nine languages using several LLMs and present an in-depth analysis in this work. Our findings reveal that: (i) deactivating the all-shared neurons significantly decreases performance; (ii) the shared neurons play a vital role in generating responses, especially for the all-shared neurons; (iii) neuron activation patterns are highly sensitive and vary across tasks, LLMs, and languages. These findings shed light on the internal workings of multilingual LLMs and pave the way for future research. We release the code to foster research in this area. 111https://github.com/weixuan-wang123/multilingual-neurons
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in recent studies, excelling in both understanding and generating text across various languages (OpenAI, 2023; Zhang et al., 2023; Zhao et al., 2024a). Despite their proven effectiveness, the intricate mechanisms underlying their processing remain largely opaque. This opacity has given rise to a growing field of research aimed at interpreting the internal workings of the Transformer architecture (Elhage et al., 2021; Yu et al., 2023). To enhance interpretability and investigate specific aspects of model behavior, researchers have increasingly focused on the components of these models. Recent studies have explored the role of Feed-Forward Networks (FFNs) within LLMs, proposing that these components function as key-value memories for storing factual and linguistic knowledge (Geva et al., 2020, 2022; Ferrando et al., 2023). While these studies have analyzed neuron behaviors based on activation states in monolingual settings, there remains a significant gap in our understanding of how neurons behave in multilingual contexts.
@@
To address this research gap, recent research attempts to unveil the mechanistic interpretability of multilingual LLMs. Bhattacharya and Bojar (2023) categorized neurons into two coarse-grained groups: language-agnostic (shared across languages) and language-specific (unique to a language). However this categorization oversimplifies the complexity observed in cross-lingual studies, where neuron overlap varies significantly between languages (Stanczak et al., 2022; Zhao et al., 2023; Liu et al., 2024). Additionally, most research has been confined to single-task analyses, overlooking how neuron types might shift across diverse tasks (Bhattacharya and Bojar, 2023; Tang et al., 2024; Tan et al., 2024). This underscores the need for a more nuanced, fine-grained classification method to enhance our understanding of the multifaceted roles of neurons in multilingual LLMs.
@@
In this work, our research introduces a fine-grained classification of neurons, enabling a detailed exploration of their functions across languages. For a specific English example and its translations in eight other languages, we categorize neurons into four distinctive types (see Figure 1): all-shared neurons, which remain active for all the inputs regardless of language; partial-shared neurons, which are activated only for inputs in certain languages; specific neurons, which are activated exclusively for inputs in one language; and non-activated neurons, which are not activated for any inputs. We begin by analysing the importance of each neuron type by deactivating them individually. Then we probe their contributions to generating answers using the Generation Impact Score (Geva et al., 2022) and the Correctness Impact Score (Voita et al., 2023). Furthermore, by examining the percentage of neurons in each type, we analyse activation patterns to gain insights into the internal workings of LLMs. We systematically study neuron behaviours across three distinct tasks, including reasoning, fact probing, and question answering, in nine languages. This analysis utilizes diverse model backbones such as BLOOMZ-7b, LLaMA2-7b-chat, BLOOM-7b, and XGLM.
@@
We provide substantial empirical evidence detailing neuron contributions and activation patterns in this study, leading to several significant findings. Here are the main takeaways:
All-shared neurons have a significant impact on model performance. We individually deactivate each type of neurons in LLMs and observe substantial performance declines (up to 87.39%) across tasks (see Section 5).
All-shared neurons are crucial in generating responses. Both the Generation Impact Score and Correctness Impact Score highlight the significance of the shared neurons in the generation process, and the all-shared neurons make substantially more contributions compared to other neuron types (see Section 6).
@@
Neuron activation patterns vary across tasks, LLMs, and languages. We observe that the patterns of four types of neurons vary across tasks (see Section 7.2) and LLMs (see Section 7.3). Moreover, our empirical results show that languages from the same language family do not always exhibit a higher degree of neuron sharing compared with languages from distinct language families (see Section 7.4).
@@ 2 Related Work
For multilingual neuron analysis, Bhattacharya and Bojar (2023) explored the neuron sharing between two languages. Tang et al. (2024); Tan et al. (2024); Liu et al. (2024); Kojima et al. (2024) classified neurons in an FFN block to language-specific and language-agnostic based on predefined threshold. However, the broad classification into two groups is inadequate for detailed multilingual analysis. Additionally, these studies classified neurons based on the single task (Tan et al., 2024; Liu et al., 2024), without considering the potential adaptation of neurons under various languages and semantics brought forth by inputs from various multilingual tasks. We investigate neurons’ behaviors across multiple languages and tasks to this end.
3 Fine-Grained Neuron Classification
| XNLI | KE (EN ALL) | KE (ALL EN) | Fact Probing | |||||||||
| pct. | pct. | pct. | pct. | |||||||||
| baseline | 0.00% | 41.99 | 0.00% | 0.00% | 38.39 | 0.00% | 0.00% | 41.74 | 0.00% | 0.00% | 41.98 | 0.00% |
| w/o. all | 9.92% | 9.38 | -77.66% | 8.71% | 4.84 | -87.39% | 10.17% | 13.19 | -68.40% | 0.28% | 21.86 | -50.31% |
| w/o. partial | 10.33% | 42.65 | 1.57% | 13.36% | 40.67 | 5.94% | 10.55% | 39.59 | -5.15% | 36.73% | 26.86 | -36.02% |
| w/o. specific | 3.14% | 42.07 | 0.19% | 4.91% | 40.78 | 6.23% | 3.82% | 40.77 | -2.32% | 16.56% | 12.68 | -67.41% |
| w/o. non-act. | 76.61% | 35.90 | -14.50% | 73.22% | 21.96 | -42.80% | 75.46% | 19.58 | -53.09% | 46.43% | 26.68 | -36.45% |
| w/o. random | 5.00% | 42.30 | 0.74% | 5.00% | 30.98 | -19.30% | 5.00% | 41.29 | -1.08% | 1.00% | 37.86 | -9.81% |
| 15.00% | 43.13 | 2.71% | 15.00% | 31.74 | -17.32% | 15.00% | 42.14 | 0.96% | 15.00% | 35.38 | -15.72% | |
| 25.00% | 43.98 | 4.74% | 25.00% | 32.40 | -15.60% | 25.00% | 42.28 | 1.29% | 35.00% | 41.78 | -0.48% | |
| 75.00% | 36.58 | -12.88% | 75.00% | 13.29 | -65.38% | 75.00% | 16.50 | -60.47% | 45.00% | 17.06 | -59.36% | |
In this section, we provide a detailed description of the 4-way neuron classification that we propose. We begin with some background concerning neurons in the FFN block (Section 3.1). Following this, we define the four types of neurons (Section 3.2).
@@ 3.1 Neurons in FFN Blocks
A neuron inside the FFNs is defined as a linear transformation of an input representation followed by a non-linear activation (Tang et al., 2024). Every FFN block at layer involves two linear transformations separated by a point-wise activation function. Biases are omitted for brevity:
where , are linear parameter matrices, and is a non-linear activation function, where rows in and columns in are viewed as -dimensional keys and values , respectively. is the count of neurons. And the output of neurons determines the weighting of the corresponding values in .
@@
Following Voita et al. (2023); Bhattacharya and Bojar (2023); Tang et al. (2024), we define a neuron as activated when its activation value satisfies . Conversely, if the activation value is , the neuron is considered deactivated.
3.2 Definitions of Four Types of Neurons
In this work, we categorize the neurons into four types based on their activation values and detail the neuron classification in this section. To ablate the impact of semantic discrepancies across languages, the datasets used in this work are initially in English and then translated into foreign languages (see Section 4.1), so we can formulate the -th example as , where indicates the -th language and is the total number of languages. Given the -th example , the set of all-shared neurons at the -th layer can be defined as:
where is the set of all the neurons at the -th layer and is the -th neuron in . Furthermore, the non-activated neurons is the set of neurons whose activation value is less than or equal to zero in all languages, as follows:
@@ 3.2 Definitions of Four Types of Neurons
Note that, we only examine the activation state of the last token of the input, as that is when the LLM performs the prediction task.
@@ 4.1 Multilingual Tasks
We perform analysis on neurons in FFN blocks of various LLMs, harnessing their multilingual capabilities in three diverse tasks which consist of multilingual parallel sentences, including XNLI (Conneau et al., 2018), Fact Probing (Fierro and Søgaard, 2022), and Cross-lingual Knowledge Editing (KE) (Wang et al., 2023). For the Cross-lingual KE, we analyse the LLMs in two setups, including EN (Edit) ALL (Test) and ALL (Edit) EN (Test). These test sets across languages are translated from the original English test set. More details are described in Appendix C.
@@
These tasks cover nine diverse languages, including English (en), German (de), Spanish (es), French (fr), Russian (ru), Thai (th), Turkish (tr), Vietnamese (vi), and Chinese (zh). Prompts are detailed in Appendix D.
4.2 Model Backbones
@@
We mainly analyse the contributions and activation patterns of neurons in an instruction-finetuned multilingual model BLOOMZ-7b (Muennighoff et al., 2023). We also include the analysis of other multilingual LLMs: BLOOM-7b (Scao et al., 2022), LLAMA2-7b-chat (Touvron et al., 2023), and XGLM (Lin et al., 2022). We use one NVIDIA A100 (40G) for all experiments.
5 Shared Neurons Are Crucial to Performance